# KunServe: Efficient Parameter-centric Memory Management for LLM Serving

### v4

@@
Abstract
Serving LLMs with a cluster of GPUs is common nowadays, where the serving system must meet strict latency SLOs required by applications. However, the stateful nature of LLM serving requires maintaining huge states (i.e., KVCache) in limited GPU memory. Under spikes in real-world workloads, GPU memory can be easily throttled, leading to orders of magnitude higher response latency due to queuing introduced by waiting for KVCache to be reclaimed. Prior KVCache-centric approaches handle load throttling by dropping, migrating, or swapping KVCache. These methods fail to release sufficient memory quickly with requests still queued.
This paper proposes the first parameter-centric approach to handling throttling by selectively dropping replicated parameters to instantly free memory for requests, based on an unnoticed observation that model parameters are commonly replicated across GPUs for serving LLMs. With additional memory, all requests can be served with a larger batch without queuing. To make the parameter-centric approach correct and efficient, we cooperatively execute requests on GPUs with a complete copy of parameters using pipeline parallelism, and derive an appropriate drop plan without unnecessary cooperation. We also design techniques to minimize the performance overhead due to pipeline parallelism with the execution patterns of requests under drop. Evaluations show that KunServe reduces the tail TTFT of requests under throttling by up to 72.2 $\times$ compared to the state-of-the-art systems including Llumnix, vLLM and InferCept.
footnotetext: Work done during Yuxin Lai (HUST)’s internship at SJTU.
@@ Introduction
Transformer-based large language models (LLMs) are reshaping the computing industry, which generate output in a token-by-token streaming fashion. The tokens are used by downstream tasks like chatbots [33], copilots [23], and interactive agents [21]. Such tasks require human interactions, so serving LLMs has tight latency requirements, e.g., less than 1 second [11, 50]. The smaller, the better [22]. Specifically, both the time to generate the first token (TTFT) and the time between subsequent tokens (TPOT) are important metrics.
@@
A key feature of LLM inference is that the computation is stateful: before generating the final token, the intermediate results (typically termed KVCache) are kept in the scarce GPU memory (HBM). Such a stateful generation introduces a key challenge: the serving latency could spike (up to 239 $\times$ in BurstGPT [45], see §2.2 and others in §5) when the stored KVCache throttles the precious HBM. Such throttling is common under real-world request bursts [35, 20] since the KVCache is proportional to the number of requests processed (or to be processed). Such throttling significantly impacts latencies, because requests must wait for GPUs to free up sufficient memory for processing. Unfortunately, it could take seconds for LLMs to release memory due to the long and unpredictable token generation process.
@@
State-of-the-art approaches adjust KVCache to handle memory throttling [27, 46, 37, 41]. When a GPU lacks sufficient HBM and causes request queuing, the system either drops KVCache, swaps it out, or migrates it to an available spare GPU to make room for queued requests (detailed in §2.3). We argue that adjusting KVCache does not fundamentally resolve the queuing problem during memory throttling, because these methods don’t release sufficient memory quickly, they replace one set of queued requests with another, without enabling all requests to make progress. Thus, a portion of requests must still be queued, resulting in sharp tail latency increases (e.g., more than 100 $\times$).
@@
This paper answers a key question: how can we effectively handle the latency spikes caused by memory throttling in LLM serving? To answer this question, we propose a new mechanism—parameter-centric memory management—to instantly free up abundant GPU memory upon throttling, so no requests need to be queued. The method is motivated by two insights. First, the HBM usage is dominated by both KVCache and model parameters (34–74% per GPU, see Table 1), so dropping a portion of parameters can free up sufficient memory for processing all requests. While intuitive, dropping parameters inevitably disrupts the inference process, making the GPUs with dropped parameters unable to process requests. Thus, our second insight is that, due to the massive computational requirements of model serving, modern LLMs are served with a cluster of GPUs where the parameters are replicated across multiple GPUs [41, 12, 34, 10]. As a result, as long as we carefully drop parameters to ensure complete copies exist cluster-wide, we can correctly process requests with dropped parameters.
@@
Our parameter-centric memory management works in a three-step process. First, upon detecting some GPU has suffered or is about to suffer from memory throttling, we derive a drop plan across GPUs to free up sufficient memory. Afterward, requests executed on GPUs with dropped parameters are re-scheduled to groups of GPUs with complete parameters. Requests are then executed with parallel inference techniques across GPUs with pipeline parallelism. Finally, once the memory demand of KVCache decreases, we restore parameters on the original GPUs to achieve the lowest inference latency.
@@
While it sounds simple, effectively realizing the aforementioned parameter-centric memory management requires addressing several challenges. First, generating an efficient drop plan should holistically consider the memory freed up by the dropped parameters as well as the performance overhead introduced by dropping too many parameters. Meanwhile, we need a system mechanism to allow existing GPU kernels highly optimized for LLMs to use the HBM freed up by dropped parameters without modifications. To this end, we first leverage the predictable performance pattern of pipeline parallelism—the more parameters dropped, the more performance overhead incurred—to derive a drop plan that minimizes the performance overhead while providing sufficient memory. Next, we design a unified GPU virtual memory management system with GPU’s advanced virtual memory features [4] to allow unmodified kernels to access the memory used for parameters for KVCache (§4.1).
@@
Second, efficiently resuming requests after dropping requires exchanging KVCache between GPUs, since it is coupled with the parameters. Doing so is non-trivial: exchanging KVCache through the network would cause significant interference with pipelined execution requests, because transferring large KVCache saturates the network used for forwarding activations. Observing that the activation transfer is more critical and the network usage is small, we design a coordinated network transfer engine that prioritizes the activation transfer to ensure both transfers are not interfered (§4.2).
@@
Finally, we faces the problem of pipelined bubbles [6] when executing requests across GPUs with dropped parameters inevitably, resulting in increased serving latencies and degraded throughput. Such throughput degradation is harmful particularly in our setup because if requests are processed at a slower rate, it could lead to another round of memory throttling. To tackle the problem, we identify the root cause of bubbles due to sub-optimal batch formulation in state-of-the-art systems like Sarathi-Serve [6]. By leveraging the observation that under throttling many requests are queued, we holistically form microbatches of queued requests with a new execution estimation metric combined with a lookahead batch formulation algorithm. The scheduling minimizes the pipelined bubbles (§4.3).
@@
We built KunServe, the first LLM serving system with parameter-centric memory management. Under various real-world traces and datasets, KunServe achieves up to 12.7-72.2 $\times$ tail latency decrease in real workloads and reduces 7.2–12.8% SLO violations under different SLO factors compared to state-of-the-art baselines. In summary, this paper makes the following contributions:
@@
  • A new parameter-centric memory management design for coping with memory throttling (§3).
@@
  • Evaluations confirming the benefits of our design (§5).
@@
Figure 1: An illustration of a typical LLM serving scenario: the model is deployed on different servers (a) with model parallelism (b) and prefill and decode requests are processed in a batched (c) way. exe. is abbreviation for execution.
@@
LLM basics. LLM is a transformer-based [43] deep learning model. Compared with traditional DNN, a key difference is that it executes requests in an auto-regressive pattern with a prefill and decode phase. In the prefill phase, the input is fed to the model to generate the first token of the output. The decode phase then iteratively generates the rest of the output in a token-by-token way, where each iteration takes the previously generated token as well as the prefill input as the input. The decodeWe use the term decode to refer to the execution of a single iteration in the decode phase in this paper. ends when the model generates a special end-of-sequence (EOS) token.
During LLM inference, since the same prefix of input is shared across all the iterations, the internal results (termed KVCache) are cached in the GPU memory (HBM) for acceleration. This makes the computation patterns of prefill and decode different [35, 25, 50]: the prefill is compute-bound, while the decode is memory-bound. To improve GPU utilization, modern LLM inference frameworks fuse prefill and decode requests into a single batch [6, 27].
@@
Deploying LLM instances with parallelism and replication. LLMs can be deployed on a single GPU or multiple GPUs with parallelism [29, 39, 49]. Pipeline parallelism (PP) partitions model parameters by layers, where layers belonging to the same group (called stage) are executed on the same GPU. Tensor parallelism (TP) partitions each layer, while different stages can reside on the same GPU. Parallelism comes at the cost of extra latency. For methods with high communication requirements like TP, parallelism is only applied to GPUs within the same server, because their interconnects are fast. PP on the other hand, can apply to GPUs across servers thanks to its ultra-low communication volume. However, PP suffers from bubbles [7] especially for requests with a small batch size. TP and PP can be applied together.
@@
In this paper, we term the minimal number of GPUs that have a single copy of the model parameters as a serving instance. The GPUs of an instance can be within the same server or across servers, but typically within the same server for the lowest serving latency unless the model is out of the capacity of a single server, which is rare (e.g., Llama-3-405B). There are typically multiple instances with replicated models, as shown in Figure 1, because a single instance has limited serving capacity.
Problem: TTFT spikes caused by memory throttling
@@
Figure 2: Analysis of TTFT increases due to GPU memory throttling (abbreviated as “Throt.” in figure). (a) The incoming request rate of BurstGPT trace [45]. (b) KVCache memory demand on vLLM [27] and (c)–(e) requests TTFT of existing solutions (§2.3).
@@
Huge HBM demands and memory throttling of LLM serving. The overall memory demand for processing LLM is huge. Considering serving a Qwen-2.5-14B model, where each token consumes 192 KB memory, an already small amount due to GQA [8]. Nevertheless, a typical burst introduces an accumulation of 243 K tokens per GPU on BurstGPT trace (see Figure 2), consuming 45 GB KVCache memory per GPU.
@@
Given the huge memory demand of LLM serving, GPUs may encounter memory throttling for two reasons. First, real-world traces exhibit spiked loads: Figure 2 (a) shows a real-world trace on BurstGPT [45], where the incoming request rate increases by 2 $\times$ at time 45s with no clear pattern. Since the KVCache demand is also proportional to the request rate, the memory demand can easily exceed the GPU memory capacity. Second, each request’s KVCache may reside in GPU for a long time, whose time is unpredictable, e.g., depending on how long LLMs generate the EOS. For BurstGPT dataset, the average stay time for a request is 11 seconds, with a variance of 14.9 seconds. Thus, even the HBM is sufficient to hold new coming requests, GPUs still suffer from memory throttling due to the unfinished request.
@@
Figure 2 (b) shows how existing serving system behaves under BurstGPT. During a 640s serving period (§5.4), we observed two throttling events on vLLM [27], a state-of-the-art LLM serving system. The timing of throttling is strongly related to the request spikes. Note that we have chosen a practical setup where the overall HBM provisioned for KVCache is 2.1 $\times$ higher than the average requirement. We use a standard approach [41] that counts the memory demands by considering both the in-processing requests and head-of-line queuing requests.
@@
TTFT spikes. GPU memory throttling is a killer for the serving performance. As shown in Figure 2 (c), the TTFT increases significantly after the throttling happens (see (b)). The increase comes from the queuing delays of waiting for sufficient memory to be freed up. The queuing time can be lengthy because the memory can only be freed once the ongoing request batch finishes. As we have mentioned before, the ongoing requests may take a long time to finish (e.g., up to 150s in BurstGPT).
@@
Figure 3: Existing methodologies (a)—(c) to address memory throttling of KVCache and how KunServe tackles this issue (d) via parameter dropping (❶) and remapping memory to enlarge KVCache region (❷).
Drop the KVCache [27, 46, 37] (Figure 3a). A naive solution is to drop some KVCache of ongoing requests (❶). Subsequently, queued requests can be processed with the freed GPU memory (❷). However, requests with dropped KVCache must be re-enqueued and recomputed, which also suffers the queuing overhead (❸) even without considering the recomputation cost. As a result, (c) shows that simply dropping the KVCache faces up to 239 $\times$ TTFT increases during memory throttling, even with a modest average memory load (56.3%).
Swap the KVCache [50, 27, 5, 48] (Figure 3b). A classic solution to handle memory throttling is swapping: when it happens, the system swaps out the overflowed KVCache to other storage (e.g., CPU DRAM) to free the GPU memory for execution (❶). The key problem is that as the GPU memory is still insufficient, there must be requests queued, even without considering the swapping overhead. For example, under throttling, InferCept [5] concurrently swaps out the KVCache of ongoing requests to hide the transfer overhead, but the queued requests are still waiting for ongoing requests to finish. The waiting is lengthy because the overall decode time is orders of magnitude higher than TTFT. As a result, we still observed a 92 $\times$ TTFT spike on InferCept [5] in Figure 2 (d). Worse still, the swapped-out requests (❸) further suffer high TPOT (see Figure 13).
Migrate the KVCache [41] (Figure 3c). Finally, observing that a serving cluster typically has multiple instances, a recent work (Llumnix [41]) migrates requests from a memory-throttled GPU to other (relatively) spare GPUs (❶) for pending requests (❷). The observation is that while no single GPU can hold all the pending requests, we can migrate requests to reduce fragmentation to free up sufficient memory. However, the queued requests can still be stalled because memory is occupied by migrating requests or the destination node is also memory-throttled (❸). Worse still, under spike workloads, there is little room for using migration to free up memory because the overall memory KVCache is insufficient even without considering fragmentation. Thus, as shown in Figure 3 (e), migration still leads to a 148 $\times$ P99 TTFT increase (compared to the P50).
System Overview
@@
Table 1: Common LLM models and their GPU memory usage.
Approach: online parameter dropping. As mentioned in the introduction, KunServe is based on two key observations of LLM serving: (1) parameters typically take up a considerable portion of HBM per GPU (see Table 1) that can be used for KVCache and (2) parameters are replicated so dropping them for KVCache does not impact LLM serving. Figure 3 (d) illustrates KunServe’s main approach and a comparison with other baselines. When the HBM used for KVCache is exhausted on GPU0 and GPU1, it instantly drops the second half of layers on GPU0 and first half of layers on GPU1 (❶). Then, the queued requests are rescheduled on both GPUs (❷) for execution via pipeline parallelism.
Discussion: why pipeline parallelism? We chose pipeline parallelism because the network requirement can be easily satisfied with the interconnects between instances. Specifically, it requires orders of magnitude smaller communications than other parallelism setups that support execution after the parameter drop like tensor parallelism. While instances could link together via fast interconnects like NVLink for tensor parallelism, the domain of NVLink is much smaller than networks that could serve pipeline parallelism well like RDMA [32]. Thus, under throttling, we may be unable to find sufficient instances linked by NVLink.
System architecture. Figure 4 illustrates our system architecture as well as the workflow of parameter-centric memory management for handling memory throttling. KunServe is a cluster-serving system that manages a set of LLM serving instances. Requests are routed through a global dispatcher, which enqueues them to the local executor of each instance for execution. Our dispatcher incorporates the load-balancing design from Llumnix [41]. The global monitor collects usage information and calculates the load metric for each instance.
Once a memory throttling event is detected by the monitor, it invokes our global memory manager (➀) to generate dropping plans. The plan is then forwarded to the local manager on the involved instances (➁) to adjust the memory according to the plan (details in §4.1).
Figure 4: System overview of KunServe.
@@
Figure 5: A comparison of the latency of different parallelism setups on BurstGPT dataset. All setups are evaluated with 8 GPUs.
Parameter drop during memory throttling
@@
Upon throttling, KunServe needs to generate a drop plan to free up sufficient memory. Besides the memory requirement, the plan has to meet the following requirements: (1) we need to generate the plan quickly online, (2) the plan needs to ensure a correct execution and (3) the plan needs to minimize the performance loss caused by parameters drop.
@@
For (2), we only need to ensure that all the instances combined have a complete copy of parameters. However, dropping too many parameters has a performance cost. For example, suppose we are serving a 7-layer model with 7 instances. While dropping 6 layers on all instances can free 85 % of the HBM for KVCache, it forces the scheduler to split the batch into microbatches with smaller sizes, which reduces batch execution efficiency as well as are more vulnerable to pipeline bubbles. Figure 5 compares the serving latencies with the different degrees of parameter drop. We can clearly see that the more parameters dropped, the higher the execution latency.
Figure 6: The pseudocode of drop plan generation algorithm.

### v5

# KunServe: Parameter-centric Memory Management for Efficient Memory Overloading Handling in LLM Serving

@@
Abstract.
Serving LLMs with a cluster of GPUs is common nowadays, where the serving system must meet strict latency SLOs required by applications. However, the stateful nature of LLM serving requires maintaining huge states (i.e., KVCache) in limited GPU memory. Under spikes in real-world workloads, GPU memory can be easily overloaded, leading to orders of magnitude higher response latency due to queuing introduced by waiting for KVCache to be reclaimed. Prior KVCache-centric approaches handle overloading by dropping, migrating, or swapping KVCache. These methods fail to release sufficient memory quickly with requests still queued.
This paper proposes the first parameter-centric approach to handling overloading by selectively dropping replicated parameters to instantly free memory for requests, based on an unnoticed observation that model parameters are commonly replicated across GPUs for serving LLMs. With additional memory, all requests can be served with a larger batch without queuing. To make the parameter-centric approach correct and efficient, we cooperatively execute requests on GPUs with a complete copy of parameters using pipeline parallelism, and derive an appropriate drop plan without unnecessary cooperation. We also design techniques to minimize the performance overhead due to pipeline parallelism with the execution patterns of requests under drop. Evaluations show that KunServe reduces the tail TTFT of requests under overloading by up to 72.2 $\times$ compared to the state-of-the-art systems including Llumnix, vLLM and InferCept.
LLM Serving; Cloud computing; Parameter-centric memory management
[…]
@@ Introduction
Transformer-based large language models (LLMs) are reshaping the computing industry, which generate output in a token-by-token streaming fashion with auto-regressive inference. The tokens are used by downstream tasks such as chatbots (chatgpt, ), copilots (copilot, ), and interactive agents (DBLP:conf/iclr/FurutaLNMFGG24, ). Such tasks require human interaction, so serving LLMs has tight latency requirements, e.g., less than 1 second (latency-study, ; distserve, ). The smaller, the better (aws-latency, ). Specifically, both the time to generate the first token (TTFT) and the time between subsequent tokens (TPOT) are important metrics.
@@
A key feature of LLM inference is that the computation is stateful: before generating the final token, the intermediate results of previously generated tokens (termed KVCache) are kept in the scarce GPU memory (HBM) to accelerate future token generation. Such a stateful generation introduces a key issue: the serving latency could spike (up to 239 $\times$ in BurstGPT (burstgpt, ), see §2.2 and others in §5) when the stored KVCache exhausts the precious HBM. Such overloading is common under real-world request bursts (splitwise, ; DBLP:conf/osdi/FuXHBUPM24, ) since the KVCache is proportional to the number of requests processed (or to be processed). Such overloading significantly impacts latency, because requests must wait for GPUs to free up sufficient memory for processing. Unfortunately, it could take seconds for LLMs to generate the final token so as to release memory due to the long and unpredictable token generation process.
@@
State-of-the-art approaches adjust KVCache stored in GPU memory to handle overloading (vllm, ; DBLP:journals/corr/abs-2404-09526, ; DBLP:journals/corr/abs-2407-00079, ; llumnix, ). When a GPU lacks sufficient HBM and causes request queuing, the system either drops KVCache of existing requests, swaps it out, or migrates it to an available spare GPU to make room for queued requests (detailed in §2.3). We argue that adjusting KVCache does not fundamentally resolve the queuing issue caused by memory overloading, because these methods do not release sufficient memory for all requests, i.e., they replace one set of queued requests with another. Thus, a portion of requests must still be queued, still resulting in sharp tail latency increases (e.g., more than 100 $\times$).
@@
This paper answers a key question: how can we effectively handle the latency spikes caused by memory overloading in LLM serving? To answer this question, we propose a new system mechanism—parameter-centric memory management—to instantly free up abundant GPU memory upon overloading for all requests to eliminate queuing. Our method is motivated by two insights. First, the HBM usage is dominated by both KVCache and model parameters (34–74% per GPU, see Table 1), so dropping a portion of parameters can free up sufficient memory for processing all requests. While intuitive, dropping parameters inevitably disrupts the inference process, making the GPUs with dropped parameters unable to process requests. Thus, our second insight is that, due to the massive computational requirements of model serving, modern LLMs are served with a cluster of GPUs where the parameters are replicated across multiple GPUs (bedrock, ; openai-api, ; serverless-ray, ; basten-deepseek, ; preplaxity-deepseek, ; llumnix, ; splitwise, ). As a result, as long as we carefully drop parameters to ensure complete copies exist cluster-wide, we can correctly process requests with dropped parameters using cooperative execution.
@@
Our parameter-centric memory management operates in a three-step process. First, upon detecting that the serving system has suffered or is about to suffer from memory overload, we derive a drop plan and execute it across GPUs to free up sufficient memory. Afterward, requests executed on GPUs with dropped parameters are seamlessly rescheduled to groups of GPUs with complete parameters to ensure complete execution. These requests are executed using parallel inference techniques across GPUs with pipeline parallelism, since other techniques like tensor parallelism have more stringent network requirements. Finally, once the memory demand of the KVCache decreases, we restore parameters on the original GPUs and reschedule the requests accordingly to achieve the lowest inference latency.
@@
Although the idea may appear simple, achieving parameter-centric memory management necessitates tackling a set of challenges. First, generating an efficient drop plan should holistically consider the memory freed up by the dropped parameters as well as the performance overhead introduced by dropping too many parameters. Meanwhile, we need a system mechanism to allow existing GPU kernels highly optimized for LLMs to use the HBM freed up by dropped parameters without modifications. To this end, we first leverage the predictable performance pattern of pipeline parallelism—the more parameters dropped, the more performance overhead incurred—to quickly derive a drop plan that minimizes the performance overhead while providing sufficient memory. Next, we design a unified GPU virtual memory management system with advanced GPU virtual memory features (cuda-memory, ) to allow unmodified kernels to access the memory used for parameters for KVCache (§4.1).
@@
Second, efficiently resuming requests after dropping requires exchanging KVCache between GPUs, since it is coupled with the parameters. However, such an exchange would significantly interfere with the pipeline-executed requests, because transferring large KVCache saturates the network used for forwarding activations. Observing that the activation transfer is more critical and the network usage is small, we design a coordinated network transfer engine that prioritizes the activation transfer to ensure both transfers are not affected (§4.2).
@@
Finally, the pipelined execution across multiple GPUs after parameter dropping causes GPU bubbles (sarathi, ), resulting in increased serving latencies and degraded throughput. The throughput degradation is particularly harmful in our setup, because if requests are processed at a slower rate, it could lead to another round of memory overloading. To tackle this problem, we identify the root cause of bubbles as sub-optimal batch formulation in state-of-the-art systems like Sarathi-Serve (sarathi, ). By leveraging the observation that under overloading many requests are queued, we holistically form microbatches of queued requests using a new execution estimation metric combined with a lookahead batch formulation algorithm. Our scheduling minimizes the pipeline bubbles thanks to the holistic formulation during pipelined execution (§4.3).
@@
We built KunServe, the first LLM serving system with parameter-centric memory management. Under various real-world traces and datasets, when compared with the state-of-the-art baselines including Llumnix (llumnix, ), vLLM (vllm, ) and InferCept (infercept, ), KunServe achieves up to 12.7-72.2 $\times$ tail latency reduction in these workloads, which further results in 7.2–12.8% lower SLO violations under common SLO factors. In summary, this paper makes the following contributions:
@@
  • A new parameter-centric memory management design for coping with memory overloading under LLM serving (§3).
@@
  • Extensive evaluations confirming the benefits of KunServe5).
KunServe is open-sourced at https://github.com/SJTU-IPADS/kunserve.
@@
Figure 1. An illustration of a typical LLM serving scenario: (a) the model is deployed on different servers with model parallelism and prefill and decode requests are processed in a batched way. exe. is abbreviation for execution.
@@
LLM basics.    LLM is a transformer-based (DBLP:conf/nips/VaswaniSPUJGKP17, ) deep learning model. Compared with traditional DNN, a key difference is that it executes requests in an auto-regressive pattern with a prefill and decode phase. In the prefill phase, the input is fed to the model to generate the first token of the output. The decode phase then iteratively generates the rest of the output in a token-by-token way, where each iteration takes the previously generated token as well as the prefill input as the context. The decode [^1] We use the term decode to refer to the execution of a single iteration in the decode phase in this paper. ends when the model generates a special end-of-sequence (EOS) token.
During LLM inference, since the same prefix of input is shared across all the iterations, the internal results (termed KVCache) are cached in the GPU memory (HBM) for acceleration. This makes the computation patterns of prefill and decode different (splitwise, ; DBLP:journals/corr/abs-2401-11181, ; distserve, ): the prefill is compute-bound, while the decode is memory-bound. To improve GPU utilization, modern LLM inference frameworks fuse prefill and decode requests into a single batch (sarathi, ; vllm, ).
@@
Deploying LLM instances with parallelism and replication.    LLMs can be deployed on a single GPU or multiple GPUs with parallelism (DBLP:conf/osdi/LiZZL00HCZGS23, ; DBLP:journals/corr/abs-1909-08053, ; DBLP:conf/osdi/ZhengLZZCHWXZXG22, ). Pipeline parallelism (PP) partitions model parameters by layers, where layers belonging to the same group (i.e., stage) are executed on the same GPU. Tensor parallelism (TP) partitions each layer, while different stages can reside on the same GPU. Parallelism comes at the cost of extra latency. For methods with high communication requirements like TP, parallelism is only applied to GPUs within the same server, because their interconnects are fast. PP on the other hand, can apply to GPUs across servers thanks to its ultra-low communication volume. However, PP suffers from bubbles (sarathi-v1, ) especially for requests with a small batch size. TP and PP can be applied together.
@@
In this paper, we define the minimal set of GPUs that have a single copy of the model parameters as a serving instance. The GPUs of an instance can be within the same server or across servers, but typically within the same server for the lowest serving latency unless the model exceeds capacity of a single server, which is rare (e.g., Llama-3-405B). For a serving cluster, a common practice is to deploy multiple instances with replicated models (basten-deepseek, ; preplaxity-deepseek, ; llumnix, ; splitwise, ), as shown in Figure 1, because a single instance has limited serving capacity.
TTFT spikes caused by memory overloading
@@
Figure 2. Analysis of TTFT increases due to GPU memory overloading (abbreviated as “Over.” in figure). (a) The incoming request rate of BurstGPT trace (burstgpt, ). (b) KVCache memory demand on vLLM (vllm, ) and (c)–(e) requests TTFT of existing solutions (§2.3).
@@
Huge HBM demands and memory overloading of LLM serving.    The overall memory demand for LLM serving is huge. For example, when serving a Qwen-2.5-14B model, each token consumes 192 KB of memory, which is already relatively small due to the use of GQA (DBLP:conf/emnlp/AinslieLJZLS23, ), a memory-efficient attention mechanism. A typical burst still introduces an accumulation of 243 K tokens per GPU on BurstGPT trace (see Figure 2), consuming 45 GB KVCache memory per GPU.
@@
We attribute GPU memory overloading to two causes. First, real-world traces exhibit spiked loads: Figure 2 (a) shows a real-world trace on BurstGPT (burstgpt, ), where the incoming request rate increases by 2 $\times$ at time 45s with no clear pattern. Since the KVCache demand is also proportional to the request rate, the memory demand can easily exceed the GPU memory capacity. Second, each request’s KVCache may reside in GPU for a long time, with an unpredictable duration, depending on how long LLMs generate the EOS. For BurstGPT dataset, the average stay time for a request is 11 seconds, with a variance of 14.9 seconds. Thus, even the HBM is sufficient to hold incoming requests, GPUs still suffer from memory overloading due to the unfinished requests.
@@
Figure 2 (b) shows how existing serving systems behave under BurstGPT. During a 640s serving period (§5.5), we observed two overloading events on vLLM (vllm, ), a state-of-the-art LLM serving system. The timing of overloading is strongly related to the request spikes. Note that we have chosen a practical setup where the overall HBM provisioned for KVCache is 2.1 $\times$ higher than the average requirement. We use a standard approach (llumnix, ) that counts the memory demands by considering both the in-processing requests and head-of-line queuing requests.
@@
TTFT spikes.    GPU memory overloading severely degrades serving performance. As shown in Figure 2 (c), the TTFT increases significantly after the overloading happens (see (b)). The increase comes from the queuing delays while waiting for sufficient memory to be freed. The queuing time can be lengthy because the memory can only be freed once the ongoing request batch finishes. As we have mentioned before, the ongoing requests may take a long time to finish (e.g., up to 150s in BurstGPT).
@@
Figure 3. (a)—(c) Existing methodologies to address memory overloading of KVCache. (d) How KunServe tackles this issue via parameter dropping (❶) and remapping memory to enlarge KVCache region (❷).
Drop the KVCache (vllm, ; DBLP:journals/corr/abs-2404-09526, ; DBLP:journals/corr/abs-2407-00079, ) (Figure 3 (a)).    A straightforward solution is to drop some KVCache of ongoing requests (❶). Subsequently, queued requests can be processed with the freed GPU memory (❷). However, requests with dropped KVCache must be re-enqueued and recomputed, which also suffers the queuing overhead (❸) even without considering the recomputation cost. As a result, Figure 2 (c) shows that simply dropping the KVCache faces up to 239 $\times$ TTFT increases during memory overloading, even with a modest average memory load (56.3%).
Swap the KVCache (distserve, ; vllm, ; infercept, ; fastserve, ) (Figure 3 (b)).    A classic solution to handle memory overloading is swapping: when it happens, the system swaps out the overflowed KVCache to other storage (e.g., CPU DRAM) to free the GPU memory for execution (❶). The key problem is that as the GPU memory is still insufficient, there will inevitably be queued requests, even without considering the swapping overhead. For example, under overloading, InferCept (infercept, ) concurrently swaps out the KVCache of ongoing requests to hide the transfer overhead, but the queued requests are still waiting for ongoing requests to finish. The waiting time can be substantial because the overall decode time is orders of magnitude higher than TTFT. As a result, we still observed a 92 $\times$ TTFT spike on InferCept (infercept, ) in Figure 2 (d). Worse still, the swapped-out requests (❸) further suffer high TPOT (see Figure 13).
@@
Qwen-3-235B 479 GB 8 (640 GB) 74.8
DeepSeek-V3-671B 1,572 GB 32 (2,560 GB) 61.4
Table 1. Popular LLM models, their parameter memory usage, the number of GPUs belonging to an instance, and the parameter memory usage ratio. Note that within an instance, Qwen-3-235B and DeepSeek-V3-671B are configured with expert parallelism with degrees 8 and 32, respectively, a common serving setup (DBLP:journals/corr/abs-2412-19437, ).
Migrate the KVCache (llumnix, ) (Figure 3 (c)).    Finally, observing that a serving cluster typically has multiple instances, a recent work (Llumnix (llumnix, )) migrates requests from a memory-overloaded GPU to other (relatively) spare GPUs (❶) for pending requests (❷). The observation is that while no single GPU can hold all the pending requests, we can migrate requests to reduce fragmentation to free up sufficient memory. However, the queued requests can still be stalled because memory is occupied by migrating requests or the destination node is also memory-overloaded (❸). Worse still, under spike workloads, there is little room for using migration to free up memory because the overall memory KVCache is insufficient even without considering fragmentation. Thus, as shown in Figure 2 (e), migration still leads to a 148 $\times$ P99 TTFT increase (compared to the P50).
System Overview
Approach: online parameter dropping.    As mentioned in the introduction, KunServe is based on two key observations of LLM serving: (1) parameters typically take up a considerable portion of HBM per GPU (see Table 1) that can be used for KVCache and (2) parameters are replicated across instances so dropping them for KVCache does not impact LLM serving. Figure 3 (d) illustrates KunServe’s main approach and a comparison with other baselines assuming two instances and each instance uses one GPU. When the HBM used for KVCache is exhausted on GPU0 and GPU1, we instantly drop the second half of layers on GPU0 and the first half of layers on GPU1 (❶). Then, the queued requests are rescheduled on both GPUs (❷) for execution via pipeline parallelism.
Discussion: why pipeline parallelism?    We chose pipeline parallelism because the network requirement can be easily satisfied with the interconnects between instances. Specifically, it requires orders of magnitude smaller communications than other parallelism setups that support execution after the parameter drop like tensor parallelism. While instances could link together via fast interconnects like NVLink for tensor parallelism, the domain of NVLink is much smaller than networks that could serve pipeline parallelism well like RDMA (nvidia-compute-network, ). Thus, under overloading, we may be unable to find sufficient instances connected by NVLink.
System architecture.    Figure 4 illustrates our system architecture as well as the workflow of parameter-centric memory management for handling memory overloading. KunServe is a cluster-serving system that manages a set of LLM serving instances. Requests are routed through a global dispatcher, which enqueues them to the local executor of each instance for execution. Our dispatcher incorporates the load-balancing design from Llumnix (llumnix, ). The global monitor collects usage information and calculates the load metric for each instance.
Once a memory overloading event is detected by the monitor, it invokes our global memory manager (➀) to generate dropping plans. The plan is then forwarded to the local manager on the involved instances (➁) to adjust the memory according to the plan (details in §4.1).
@@
Figure 5. A comparison of the latency of different parallelism on BurstGPT dataset. All setups are evaluated with 8 GPUs.
Parameter drop during memory overloading
@@
Upon overloading, KunServe needs to generate a drop plan to free up sufficient memory. Besides the memory requirement, the plan has to meet the following requirements: (1) we need to generate the plan quickly online, (2) the plan needs to ensure a correct execution and (3) the plan needs to minimize the performance loss caused by parameters drop.
@@
For (2), we only need to ensure that all the instances combined have a complete copy of parameters. However, dropping too many parameters incurs a performance cost. For example, suppose we are serving a 7-layer model with 7 instances. While dropping 6 layers on all instances can free 85 % of the HBM for KVCache, it forces the scheduler to split the batch into microbatches with smaller sizes, reducing the GPU batch execution efficiency (flashinfer, ) and making the system more vulnerable to pipeline bubbles. Figure 5 compares the serving latencies for different degrees of parameter dropping. We can clearly see that the more parameters are dropped, the higher the execution latency.
Figure 6. The pseudocode of drop plan generation algorithm.